OpenStudy (anonymous):

[6.03] What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 12 -2x + 3y = 6 -14 24 14 -24

OpenStudy (anonymous):

[6.03] What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 12 -2x + 3y = 6 -14 24 14 -24

OpenStudy (anonymous):

how do i set up an equation for this??

OpenStudy (harsimran_hs4):

well it`s easy 1. eliminate one variable method : either find x in terms of y or y in terms of x from both the equations and equate them to get x or y probably find x in terms of y from both the equations and equate them to get the value of y

OpenStudy (anonymous):

so my answer is 14?? if i did it right..

OpenStudy (harsimran_hs4):

-14 is what i get

OpenStudy (anonymous):

this may be asking to much but how did u set up your problem??

OpenStudy (harsimran_hs4):

x = (12 + 6y)/3 from 1st and x = (6 - 3y) /-2 from 2nd so solving for y from both by comparing x gives us y

OpenStudy (anonymous):

get the quation of x from 1st equation then put that in 2nd equation.

OpenStudy (anonymous):

thank you so much

OpenStudy (harsimran_hs4):

:)

OpenStudy (anonymous):

Which point lies in the solution set for the following system of inequalities? y > x + 4 y > -x + 2 (1, 6) (-2, -1) (-3, 2) (1, 1) ..i got 1,6 a i correct??

OpenStudy (anonymous):

am i

OpenStudy (anonymous):

Yes, you are. 1,6 satisfies both inequalities at the same time.

OpenStudy (anonymous):

Create a unique example of dividing a polynomial by a monomial and provide the simplified form. Explain, in complete sentences, the two ways used to simplify this expression and how you would check your quotient for accuracy.

OpenStudy (anonymous):

help plz